OpenStudy (anonymous):

is it same p(a or b) = p(a)+p(b)-p(a and b) or p(a and b ) = p(a)+p(b) -p(a or b) i need to know that both are same and can we use them according to question

OpenStudy (mathstudent55):

The first one is used to find the probability of two events when they are not mutually exclusive.

OpenStudy (anonymous):

what about second

OpenStudy (mathstudent55):

The second equation is not valid.

OpenStudy (freckles):

why is the second one not valid ?

OpenStudy (freckles):

x=A+B-y is the same as x+y=A+B which is the same as y=A+B-x